# Multi-LF: A Unified Continuous Learning Framework for Real-Time DDoS Detection in Multi-Environment Networks

### v1

@@ III-A Phase 1: Data Collection
We construct a comprehensive dataset that includes both IoT and traditional (non-IoT) traffic, encompassing benign and malicious categories. To achieve this, we employ a two-pronged approach. First, we leverage the DDoShield-IoT testbed [5], which integrates NS-3 (a discrete-event network simulator) with Docker containers. This combination provides a highly realistic environment where actual binaries run in containers and communicate over a simulated network, enabling the generation and capture of real-world network traffic. We use DDoShield-IoT to generate the two types of traffic it natively supports—benign traditional traffic and malicious IoT traffic. Second, to incorporate the other two required traffic categories (benign IoT traffic and malicious traditional traffic), we integrate external datasets published in the literature. Below, we discuss these two steps in detail, and then we discuss how we merge these different traffic sources in a unified manner.
@@
    Traffic Generation Using DDoShield-IoT: DDoShield-IoT natively supports the generation of certain traffic types. Specifically, it facilitates the creation of benign traditional traffic and malicious IoT traffic. For the benign traditional traffic, it generates three types of traffic, which include FTP, HTTP, and RTMP-based video streams from legitimate services within Docker containers. For malicious IoT traffic, we leverage the Mirai malware binaries hosted in the Docker environment, in which we produce IoT-based DDoS attack traffic. In particular, we consider three classic Mirai-based attack types: ACK flood, SYN flood, and UDP flood, which reflect key threats faced by IoT devices in modern network environments.
@@
    Integrating External PCAP Datasets: While DDoShield-IoT allows us to generate both benign traditional traffic and malicious IoT traffic, it does not natively support benign IoT traffic or malicious traditional traffic. To address these gaps, we incorporate external PCAP datasets that have been widely used and validated in other research works. This approach ensures our final dataset is both comprehensive and realistic.
    For the benign IoT traffic, we use the MQTT-IoT-IDS2020 dataset [^1] https://paperswithcode.com/dataset/mqtt-iot-ids2020. The MQTT-IoT-IDS2020 dataset contains benign IoT network traces, focusing on IoT devices and services employing the MQTT protocol. It includes normal operations of IoT devices in a controlled environment, providing a clear baseline of legitimate IoT behavior. By integrating this dataset, we incorporate authentic IoT benign traffic that closely represents everyday device activities.
    For the malicious traditional traffic, we use the CIC DDoS 2019 dataset [^2] https://www.unb.ca/cic/datasets/ddos-2019.html. The CIC DDoS 2019 dataset includes a variety of DDoS attacks captured in a realistic testbed environment. These attacks are carried out on traditional (non-IoT) services and include multiple DDoS vectors that target typical enterprise or data center hosts. Incorporating this dataset enhances the representativeness of our overall traffic corpus, ensuring that our MTDS can detect malicious patterns in both IoT and traditional network segments.
To combine these different traffic sources, we employ a Python script that is running in a Docker container node, which is running in promiscuous mode within the DDoShield-IoT simulation. Our Python script receives both the live traffic generated by the DDoShield-IoT environment and batches of packets randomly selected from the external PCAP datasets. The Python script processes incoming packets without distinction between their origins.
@@
To ensure that traffic originating from different sources—DDoShield-IoT and external PCAP files—can be meaningfully compared, we apply a uniform feature extraction and aggregation methodology. First, all packets are processed through the same feature extraction pipeline (detailed in III-B), which derives both basic packet-level features (e.g., IP addresses, port numbers, protocol types, and payload sizes) and advanced statistical features computed over fixed time intervals (e.g., the average packet size, the packet rate, and the distribution of protocols observed within a given time window).
@@
This uniformity means that regardless of whether a packet is captured from the simulated environment or sourced from an external dataset, the same set of features is extracted using identical definitions, thresholds, and time windows. By treating all traffic in this consistent manner, we eliminate biases that could arise from applying different techniques or settings to different portions of the data. As a result, the final combined dataset presents a coherent feature space, making it possible to directly compare, combine, and evaluate data from multiple origins using a single, unified analytic framework.
III-B Phase 2: Feature Extraction
@@
After capturing traffic into PCAP files, we extracted features for training ML models. These features consist of two types: general features and statistical features. General features are directly related to traffic characteristics, while statistical features are derived from analyzing general features over time to detect patterns indicative of DDoS attacks.
@@
The extracted general features from the network traffic data are essential for comprehensive analysis and detection of malicious activities. The timestamp records the precise time of packet capture, allowing for chronological analysis. Source and destination IP addresses (ip_src, ip_dst) identify the communication endpoints, crucial for tracing attack origins and targets. The protocol field specifies the transport protocol (e.g., TCP, UDP), aiding in protocol-specific analysis. Source and destination ports (src_port, dst_port) indicate the application-level endpoints, useful for identifying targeted services. The presence of TCP and UDP flags (tcp_flag, udp_flag) denotes the protocol used, while Time to Live (TTL) reveals the packet’s hop count, indicating network topology and potential anomalies. Flags such as ACK, SYN, FIN, PSH, URG, and RST in TCP packets provide insights into the connection states and potential malicious behaviors like SYN flooding or connection resets. Sequence and acknowledgment numbers are vital for reconstructing TCP sessions and detecting session hijacking or manipulation. Finally, packet size and payload size metrics help in identifying unusual packet structures and potential payload-based attacks. Together, these features provide a detailed view of network traffic, enabling effective identification and mitigation of diverse cyber threats.
    ’Timestamp’, ’Source’, ’Destination’,
    ’Protocol’, ’SrcPort’, ’DstPort’,
    ’TCP’, ’UDP’, ’TTL’, ’ACK’,
    ’SYN’, ’FIN’, ’PSH’, ’URG’,
    ’RST’, ’SequenceNumber’, ’PacketSize’,
@@
These general features provide a detailed description of each packet and are essential for identifying the characteristics of network traffic. However, these features alone are not sufficient to detect DDoS attacks effectively, as DDoS attacks rely on continuous high-frequency packet transmission [51]. Therefore, we also extract statistical features calculated over fixed time windows to capture traffic patterns and anomalies. These 24 statistical features are crucial for detecting DDoS attacks and are described below:
  • Packet Count: The total number of packets observed in each time window, represented as $\text{PacketCount}_{t}=\sum_{i=1}^{n}\text{Packet}_{i}$.
  • Destination Port Entropy: Measures the entropy of destination ports to detect scanning activities, defined as $\text{Entropy}=-\sum_{i=1}^{n}p_{i}\log(p_{i})$.
  • Most Frequent Source Port: Identifies the most common source port in a window, $\text{SourcePort}_{\max}$.
  • Most Frequent Destination Port: Identifies the most common destination port in a window, $\text{DestinationPort}_{\max}$.
@@
  • Network Scanning Activity: Counts instances of SYN flags without ACK flags, $\text{Scanning}=\sum_{i=1}^{n}(\text{SYN}-\text{ACK})$.
  • Flow Rate: Calculated as packets per second, $\text{FlowRate}=\frac{\text{TotalPackets}}{\text{TimeInterval}}$.
  • Source Entropy: Entropy of source addresses, $\text{SourceEntropy}=-\sum_{i=1}^{n}q_{i}\log(q_{i})$, where $q_{i}$ is the probability distribution of source addresses.
  • Connection Errors (RST flag): Counts instances of RST flags, $\text{RSTCount}=\sum_{i=1}^{n}\text{RST}_{i}$.
  • Most Frequent Packet Size Frequency: Identifies the most common packet size, $\text{PacketSize}_{\max}$.
@@
  • Average Packet Number: Average packets per interval, $\text{AvgPackets}=\frac{\text{TotalPackets}}{\text{TimeIntervals}}$.
  • SYN Frequency: Frequency of SYN flags, $\text{SYNFrequency}=\frac{\text{TotalSYN}}{\text{TimeInterval}}$.
  • ACK Frequency: Frequency of ACK flags, $\text{ACKFrequency}=\frac{\text{TotalACK}}{\text{TimeInterval}}$.
  • TCP Frequency: Proportion of TCP packets, $\text{TCPFrequency}=\frac{\text{TotalTCP}}{\text{TotalPackets}}$.
  • UDP Frequency: Proportion of UDP packets, $\text{UDPFrequency}=\frac{\text{TotalUDP}}{\text{TotalPackets}}$.
  • Most Frequent Protocol: Most used protocol, $\text{Protocol}_{\max}$.
[…]
  • Most Frequent Payload Size: Most common payload size, $\text{PayloadSize}_{\max}$.
  • Average Payload Size: Mean payload size, $\text{AvgPayloadSize}=\frac{\text{TotalPayload}}{\text{TotalPackets}}$.
[…]
  • Average Packet Size: Mean size of packets within a window, $\text{AvgPacketSize}=\frac{\text{TotalPacketSize}}{\text{TotalPackets}}$.
In our approach, we used specific thresholds and parameters to calculate these statistical features. The processing interval was set to 1 second, which is the time interval used to aggregate packets. An abnormal size threshold was defined as 1500 bytes to identify abnormal packet sizes. The port frequency threshold was set to 5 to determine frequently used ports. Additionally, a short-lived connection threshold was defined as less than 5 packets to identify short-lived connections. These features are designed to capture various aspects of network traffic and detect deviations from normal patterns that are characteristic of DDoS attacks. By combining general features and statistical features, we can effectively distinguish between normal traffic and DDoS traffic, enhancing the detection capabilities of our system. The comprehensive feature set is illustrated in Figure 2.
Figure 2: Combining statistical and general features approach
Figure 3 shows the feature space for both general and statistical features, highlighting the importance of periodic statistical feature calculation. In this visualization approach, we use PCA to reduce the dataset’s dimensionality to three principal components and then visualize these components in a 3D scatter plot. The target labels are used to differentiate between classes, with each class represented by a unique color. It is evident that with only general features, the samples overlap, making it difficult for the model to learn distinguishable patterns. However, with statistical features, the samples are separable, aiding the model in learning distinct patterns.
Figure 3: Feature Space (a) Combined Features, (b) General Features, and (c) Statistical Features
III-C Phase 3: ML Approach
In the ML approach, we deployed several state-of-the-art methods for malicious traffic detection. We trained two models, M1 and M2. M1 is a lightweight model trained on selected samples after preprocessing while M2 is a complex model trained on a huge dataset as shown in Figure 4. This M1 model directly deals with live traffic, and if it fails, the decision is passed to M2, which then helps to learn new data where M1 failed.
Figure 4: ML models training approaches
M1: In the M1 training approach, we first selected a small sample of data by choosing 100,000 samples from each class, totaling 400,000 samples. This small dataset helps us obtain a lightweight model. Initially, we randomly selected 100,000 samples and trained different models. Then, we selected the best 100,000 samples from each class using the K-Means clustering approach [52]. K-Means clustering was applied to the standardized features, with an appropriate number of clusters chosen based on the dataset’s characteristics. For each sample, the distance to the nearest cluster center was calculated, reflecting how well each sample represented its cluster. By selecting the samples with the smallest distances to their respective cluster centers, this method ensured that the most representative samples for each class were chosen. This process was repeated for each class in the dataset, and the selected top samples from each class were compiled into a new data frame. This approach effectively reduced the dataset size while maintaining the diversity and representativeness of the original data, making it suitable for training ML models or further analysis. After selecting the important samples, we applied data scaling using the MinMax method. After that, we split the dataset with a 75% and 25% ratio, where 75% was used to train the model and 25% to test the model. We trained four models, which are described below such as stochastic gradient descent classifier (SGDC), perceptron classifier, multinomial naive Bayes (MNB), and Bernoulli naive Bayes (BNB). We used the partial_fit function to train the model in the M1 approach as we used the M1 method to retrain during continuous learning. In the end, we evaluated the performance in terms of accuracy, precision, recall, and F1 score. We also used a 10-fold cross-validation approach to evaluate the performance.
We deploy the four models in the M1 approach using their best hyperparameter settings, which were identified through a combination of literature review [11, 14, 53] and tuning within specific ranges and values using a trial-and-error approach. Table II illustrates the hyperparameter ranges and values used. Some models, such as BNB and MNB, were deployed with their default hyperparameter settings because these models are known to perform well with defaults due to their simplicity and lack of sensitivity to extensive hyperparameter tuning [54].
Below, we present a detailed discussion of each model, emphasizing their key characteristics and providing a rationale for their selection in the M-En malicious traffic detection task
TABLE II: Hyperparameter settings for the M1 models
Model Hyperparameter Value Tuning Range
SGDC loss log_loss -
max_iter 1000 [200 to 1500]
@@ III-C Phase 3: ML Approach
Perceptron max_iter 1000 [200 to 1500]
tol 1e-3 -
random_state 42 [0, 42]
MNB Default Default -
@@
III-C1 SGDC
It is a linear classifier that optimizes a linear model using stochastic gradient descent. We use it for malicious traffic detection in M-EN because it is efficient for large-scale and sparse datasets. It also supports continuous learning, which makes it suitable for our study [55]. It supports various loss functions; however, we use log loss, which helps us work with prediction probabilities to measure model confidence. The SGDC models the decision function as a linear combination of the input features as:
$z=\mathbf{w}\cdot\mathbf{x}+b=w_{1}x_{1}+w_{2}x_{2}+\ldots+w_{n}x_{n}+b$ (1)
where, $\mathbf{w}=[w_{1},w_{2},\ldots,w_{n}]$ is the weight vector, $b$ is the bias term, $\mathbf{x}$ is the input vector. Depending on the chosen loss function (log in our case), the SDGC optimizes the model using the gradient of the loss function, which can be described as:
@@ III-C1 SGDC
Then SGDC updates weights iteratively, which can be described as:
$\mathbf{w}\leftarrow\mathbf{w}-\eta\nabla L(\mathbf{w},b,\mathbf{x}_{i},y_{i})$ (3)
$b\leftarrow b-\eta\frac{\partial L(\mathbf{w},b,\mathbf{x}_{i},y_{i})}{\partial b}$ (4)
where, $\eta$ is the learning rate, $\nabla L$ is the gradient of the loss function.
III-C2 Perceptron
It is a linear classifier that makes predictions based on a simple threshold function [56]. It is used for binary classification tasks and can be extended to multiclass classification using the one-vs-rest (OvR) scheme, as in our M-En malicious traffic detection. It is also effective when data is linearly separable, which is very suitable for our case, as our dataset is highly linearly separable, as shown in Figure 3. Perceptron in sci-kit-learn also provides a continuous learning function, so we use it as M1. The Perceptron classifier aims to find a hyperplane that separates the classes. Given an input vector $\mathbf{x}=[x_{1},x_{2},\ldots,x_{n}]$, the Perceptron computes the output $y$ using:
$z=\mathbf{w}\cdot\mathbf{x}+b=w_{1}x_{1}+w_{2}x_{2}+\ldots+w_{n}x_{n}+b$ (5)
where, $\mathbf{w}=[w_{1},w_{2},\ldots,w_{n}]$ is the weight vector, $b$ is the bias term, $\mathbf{x}$ is the input vector. The output $y$ is then determined by applying the activation function:
@@ III-C2 Perceptron
III-C3 MNB
It is suitable for large datasets where fast computation is needed, as in our case [57]. It also provides the ‘partial_fit‘ function for continuous learning, which we use in our approach as the M1 model in comparison with others [58]. It estimates the probability of a class given the feature values by combining the prior probability of the class and the likelihood of the observed features given the class using Bayes’ theorem:
@@ III-C3 MNB
where, $P(C_{k}|\mathbf{x})$ is the posterior probability of class $C_{k}$, $P(C_{k})$ is the prior probability of class $C_{k}$, $P(\mathbf{x}|C_{k})$ is the likelihood of $\mathbf{x}$ given $C_{k}$, $P(\mathbf{x})$ is the marginal likelihood of $\mathbf{x}$. The likelihood is estimated based on the frequency of features:
$P(\mathbf{x}|C_{k})=\prod_{i=1}^{n}P(x_{i}|C_{k})^{x_{i}}$ (8)
where, $P(x_{i}|C_{k})$ is the probability of feature $x_{i}$ in $C_{k}$, $x_{i}$ is the count of feature $i$ in $\mathbf{x}$. Parameters are estimated using Laplace smoothing:
$P(x_{i}|C_{k})=\frac{N_{C_{k},x_{i}}+\alpha}{N_{C_{k}}+\alpha n}$ (9)
where, $N_{C_{k},x_{i}}$ is the count of $x_{i}$ in $C_{k}$, $N_{C_{k}}$ is the total count of features in $C_{k}$, $\alpha$ is the smoothing parameter. Class prediction is made by selecting the class with the highest posterior probability:
$\hat{C}=\arg\max_{C_{k}}P(C_{k})\prod_{i=1}^{n}P(x_{i}|C_{k})^{x_{i}}$ (10)
III-C4 GNB
We use it because network traffic features such as packet sizes, inter-arrival times, and other statistical metrics are continuous. GNB is well-suited for continuous data as it assumes that the features follow a Gaussian (normal) distribution [59]. It also uses Bayes’ theorem like MNB:
@@ III-C4 GNB
where, $\mu_{C_{k},i}$ is the mean of $x_{i}$ in $C_{k}$, $\sigma_{C_{k},i}^{2}$ is the variance of $x_{i}$ in $C_{k}$. Parameters are estimated using maximum likelihood estimation (MLE):
$\mu_{C_{k},i}=\frac{1}{N_{C_{k}}}\sum_{j=1}^{N_{C_{k}}}x_{j,i}$ (13)
$\sigma_{C_{k},i}^{2}=\frac{1}{N_{C_{k}}}\sum_{j=1}^{N_{C_{k}}}(x_{j,i}-\mu_{C_{k},i})^{2}$ (14)
where, $N_{C_{k}}$ is the number of samples in $C_{k}$, $x_{j,i}$ is the value of $x_{i}$ for the $j$-th sample in $C_{k}$. Class prediction is made by selecting the class with the highest posterior probability:
$\hat{C}=\arg\max_{C_{k}}P(C_{k})\prod_{i=1}^{n}P(x_{i}|C_{k})$ (15)
M2: In the M2 training approach, we trained models on the full dataset, which consists of 40 million samples. We applied data scaling and then split the dataset with a 75% and 25% ratio, where 75% was used to train the model and 25% to test the model. We trained several models such as MLP [60], LR [61], SGDC [62], BNB [63], GNB [64], random forest (RF) [65], AdaBoost (ADA) [66], support vector classifier (SVC) [67], and k-nearest neighbors (KNN) [68]. We evaluated the performance of these models in terms of accuracy, precision, recall, and F1 score. After training both M1 and M2 models, we saved the models using the pickle library and stored them with a .pkl extension to use them during real-time testing. Similar to the M1 approach, we also deploy the M2 approach models using their best hyperparameter settings, identified through a combination of literature review and tuning within specific ranges and values using an iterative experimentation approach [69, 10, 53]. Table III illustrates the hyperparameter ranges and values used. Other models, such as BNB, GNB, and MNB, were deployed with their default hyperparameter settings.
TABLE III: Hyperparameter settings for the M2 models

### v2

@@ 3.1 Phase 1: Data Collection
We construct a comprehensive dataset that encompasses both IoT and traditional (non-IoT) traffic, including both benign and malicious categories. To achieve this goal, we employ a two-pronged approach. First, we leverage DDoShield-IoT [devivo2024ddoshield], a testbed that leverages the integration of Docker with NS-3 to host actual binaries that communicate over a simulated network, mimicking a realistic network environment. This setup enables the generation and capturing of real-world network traffic. We use DDoShield-IoT to generate the two types of traffic it natively supports: benign traditional network traffic and malicious IoT network traffic. Second, to incorporate the other two missing traffic categories (benign IoT traffic and malicious traditional traffic), we integrate external PCAP datasets (published in the literature) into DDoShield-IoT. Below, we discuss these two steps in detail, and then we discuss how we merge these different traffic sources in a unified manner.
@@
    Traffic Generation Using DDoShield-IoT: DDoShield-IoT supports the generation of certain traffic types. Specifically, it facilitates the creation of benign traditional traffic and malicious IoT traffic. For the benign traditional traffic, it generates different types of traffic (e.g., HTTP/HTTPS and video traffic) from legitimate services within Docker containers. For malicious IoT traffic, we utilize the Mirai malware [mohanty2018control] (a notorious IoT botnet) binaries hosted in a Docker container in DDoShield-IoT, from which we generate IoT-based attack traffic. In particular, we generate three Mirai-based attack classes: ACK flood, SYN flood, and UDP flood, which reflect key threats by IoT botnets in modern network environments [obaidat2023creating].
@@
    Integrating External PCAP Datasets: While DDoShield-IoT allows us to generate both benign traditional traffic and malicious IoT traffic, it does not natively support benign IoT traffic or malicious traditional traffic. To address these gaps, we incorporate external PCAP datasets that have been proposed by other research works. This approach ensures our final dataset is both comprehensive and realistic.
    For the benign IoT traffic, we use the MQTTIoTIDS2020 dataset [^2] https://paperswithcode.com/dataset/mqtt-iot-ids2020. The MQTT-IoT-IDS2020 dataset contains benign IoT network traces, focusing on IoT devices and services that utilize the MQTT protocol. It includes normal operations of IoT devices in a controlled environment, providing a clear baseline of legitimate IoT behavior. By integrating this PCAP dataset, we incorporate authentic IoT benign traffic that closely represents everyday device activities.
    For the malicious traditional traffic, we use the CICDDoS2019 PCAP dataset [^3] https://www.unb.ca/cic/datasets/ddos-2019.html. This dataset includes a variety of attacks captured in a realistic testbed environment. These attacks are carried out on traditional (non-IoT) services and include multiple attack vectors that target typical enterprise or data center hosts. Incorporating this dataset enhances the representativeness of our overall traffic corpus, ensuring that our traffic detection can identify malicious patterns in both IoT and traditional network segments.
To combine these heterogeneous traffic sources, we deploy a Python sniffer running in a promiscuous Docker node inside DDoShield-IoT so it sees every frame traversing the simulated network. At each injection interval, we select a PCAP file from the external corpus and draw a random offset anywhere between its first and last packet. Rather than replaying traffic from that arbitrary byte position (which could split a session midstream), we step back to the first packet of the same flow (identified by the customary five-tuple of source IP, destination IP, source port, destination port, and transport-layer protocol [rfc3954]) and begin replaying from that clean boundary. The packets emitted from that point forward are merged with the live DDoShield-IoT traffic and passed, without source labels, into the feature-extraction pipeline (discussed in §3.2). Because both the PCAP file and the insertion point change every time, the resulting stream provides an unbiased, temporally coherent mix of IoT and traditional traffic segments.
@@
Attack Coverage: Across the two malicious sources, our M‑En dataset spans fifteen distinct attack classes. DDoShield-IoT (Mirai) contributes three IoT‑centric floods (SYN, ACK, and UDP) that reflect the malware’s most common attacks [obaidat2023creating]. The external CICDDoS2019 traces add twelve additional traditional attacks: NTP, PortScan, DNS, LDAP, MSSQL, NetBIOS, SNMP, SSDP, UDP, UDP‑Lag, WebDDoS, SYN, and TFTP floods. In addition, the MQTTIoTIDS2020 dataset provides two reconnaissance scenarios (Sparta SSH brute force and a UDP scan), which we include to capture low-volume probing behavior. This combination gives our corpus both high‑rate floods and short‑burst probing activity, providing broad coverage for M‑En detection research. Because DDoShield-IoT Docker and NS‑3 architecture lets us plug in new binaries or replay additional PCAP, future attack classes can be integrated with minimal effort, either by implementing the attack inside the testbed or by importing an annotated PCAP and following the same injection procedure described above.
Attack Coverage SYN flood, ACK flood, and UDP flood, NTP, PortScan, DNS, LDAP, MSSQL, NetBIOS, SNMP, SSDP, UDP, UDP-Lag, WebDDoS, SYN, and TFTP flood, Sparta SSH brute-force, UDP scan
3.2 Phase 2: Feature Extraction
@@
After capturing all M-En traffic, we apply a uniform feature extraction and aggregation methodology. This uniformity means that regardless of whether a packet is captured from the simulated environment or sourced from an external PCAP dataset, the same set of features is extracted using identical definitions, thresholds, and time windows. By treating all traffic in this consistent manner, we eliminate biases that could arise from applying different techniques or settings to different portions of the data. As a result, this process presents a coherent feature space, making it possible to directly compare, combine, and evaluate M-En data from multiple origins using a single, unified analytic framework.
@@
We categorize the extracted features into two types: general features (e.g., protocol types and payload sizes) and statistical features computed over fixed time intervals (e.g., the average packet size and the packet rate). General features are directly related to traffic characteristics, while statistical features are derived from analyzing general features over time to detect patterns indicative of malicious attacks.
@@
The general features are essential for a comprehensive analysis and detection of malicious activities. The timestamp records the precise time of packet capture, allowing for chronological analysis. Source and destination IP addresses (ip_src, ip_dst) identify the communication endpoints, crucial for tracing attack origins and targets. The protocol field specifies the transport protocol (e.g., TCP, UDP), aiding in protocol-specific analysis. Source and destination ports (src_port, dst_port) indicate the application-level endpoints, useful for identifying targeted services. The presence of TCP and UDP flags (tcp_flag, udp_flag) denotes the protocol used, while Time to Live (TTL) reveals the packet’s hop count, indicating network topology and potential anomalies. Flags such as ACK, SYN, FIN, PSH, URG, and RST in TCP packets provide insight into connection states and potential malicious behaviors, including SYN flooding or connection resets. Sequence and acknowledgment numbers are vital for reconstructing TCP sessions and detecting session hijacking or manipulation. Finally, the packet size and payload size metrics help identify unusual packet structures and potential payload-based attacks. Together, these features provide a detailed view of network traffic, enabling effective identification and mitigation of diverse cyber threats.
General Feature
’Timestamp’, ’Source’, ’Destination’, ’Protocol’, ’SrcPort’,
’DstPort’, ’TCP’, ’UDP’, ’TTL’, ’ACK’, ’SYN’, ’FIN’,
’PSH’, ’URG’, ’RST’, ’SequenceNumber’, ’PacketSize’,
@@
Despite the value of general features (per-packet descriptors), they alone cannot flag malicious activity reliably, because many network attacks reveal themselves through temporal patterns (sustained bursts of traffic, rapid-fire connection attempts, or repeated request sequences) that only emerge when packets are viewed in aggregate [fouladi2013frequency]. Therefore, we also extract statistical features over fixed windows to expose such patterns and anomalies. These 24 statistical features are crucial for identifying the diverse range of attacks. Our feature extraction pipeline is modular, allowing for the easy integration of additional features whenever new threats or protocols emerge. Below, we list the selected statistical features and explain the rationale behind the key thresholds; the full calculation details are provided in Appendix A.
Statistical Feature
’Packet Count’, ’Destination Port Entropy’, ’Most Frequent
Source Port’, ’Most Frequent Destination Port’,
’Short-lived Connections’, ’Repeated Connection Attempts’,
’Network Scanning Activity’, ’Flow Rate, Source Entropy’,
’Connection Errors~(RST)’, ’Most Frequent Packet
Size Frequency’, ’Abnormal Packet Size Frequency’, ’Sequence
Number Variance’, ’SYN Frequency’, ’ACK Frequency’,
’TCP Frequency’, ’UDP Frequency’, ’Packet Size Variability’,
’Average Packet Size’, ’Average Payload Size’
Thresholds & Parameters For Simulation: In our approach, we use specific thresholds and parameters to calculate these statistical features as shown in Table 2. We set the processing interval to 1 s, a value that strikes a balance between latency and statistical stability; in a preliminary sweep of 100  ms–2 s windows, 1 s produced the highest F1 score while keeping detection delay sub-second. We adopt 1500 B as the abnormal size threshold because it aligns with the default IP maximum-transmission-unit (MTU) on typical networks [rfc4821, rfc5948]; packets larger than this value ordinarily require fragmentation, so their presence is a useful indicator of atypical or crafted traffic, both of which warrant closer inspection. The port-frequency threshold is set to five occurrences per window, an empirically derived cutoff that isolates the heavy hitters responsible for a large portion of connection attempts in our training set. Finally, we label any TCP flow containing fewer than five packets as a short‑lived connection. Previous work treats such ultra‑short sessions as abnormal traffic that ends before any meaningful application‑layer exchange can occur [kotenko2017parallel, jose2021towards].
Table 2: Thresholds and Parameters Used in Simulation
Parameter Value Description
[…]
APST 1500 B Corresponds to the typical IP MTU; packets exceeding this value may indicate fragmentation or crafted traffic.
[…]
SLCT << 5 packets/flow Identifies ultra-short TCP sessions considered abnormal or incomplete exchanges.
  • PI– Processing Interval; APST– Abnormal Packet Size Threshold; PFT— Port-Frequency Threshold; SLCT— Short-Lived Connection Threshold;
These choices follow best practice in flow-analysis literature (e.g., NetFlow guidelines [rfc3954]) and are validated on a held-out subset of our dataset; nevertheless, all of these parameters remain tunable knobs in our implementation. Practitioners can re-estimate them with a grid search or Bayesian optimization to suit higher-speed links or encrypted-traffic mixes. By combining the general features with these carefully parameterized statistical features, we distinguish normal from malicious traffic more effectively.
3.3 Phase 3: ML Approach
In the ML approach, we employ several state-of-the-art methods for detecting malicious traffic. We train two models, M1 and M2. M1 is a lightweight model trained on selected samples after preprocessing, while M2 is a complex model trained on a large dataset, as shown in Figure 2. M1 directly processes live traffic, and if the confidence value of the detection is below a threshold, the decision is passed to M2, which attempts to perform the prediction (and retrain M1 based on its prediction).
Figure 2: M1 & M2 Models Training Approaches
3.3.1 Baseline M1 & M2:
In the M1 training strategy, we begin by selecting a balanced subset of the data, 100,000 samples per class, totaling 400,000 samples, to create a lightweight yet representative dataset. Initially, we train models using randomly selected samples. To refine the dataset, we then apply K-Means clustering [vzalik2008efficient] to the standardized features, selecting the top 100,000 samples per class that are closest to their respective cluster centers. This process ensures each class is represented by its most central and diverse instances. After selection, we scale the features using the MinMax method and split the dataset into 75% training and 25% testing. Then, we train four models: stochastic gradient descent classifier (SGDC), PER, multinomial naive Bayes (MNB), and Bernoulli naive Bayes (BNB). We train all models using the partial_fit function to allow continual learning. We evaluate the model’s performance using accuracy, precision, recall, and F1 score, along with 10-fold cross-validation for robustness. We select the model hyperparameters based on literature [rustam2024ai, rustam2024malicious, gada2021automated] and tune them within specific ranges and values using a trial-and-error approach. We use BNB and MNB with their default settings, which are effective due to their simplicity and low sensitivity to hyperparameters [ismail2022evaluation]. The following sections provide further analysis of each model and its suitability for the M-En malicious traffic detection task.
Table 3: Hyperparameter Settings For Baseline M1 Models
Model Hyperparameter Value Tuning Range
@@ 3.3.1 Baseline M1 & M2:
PER max_iter 1000 [200 to 1500]
@@
[…]
@@
[…]
@@
Then SGDC updates weights iteratively, which can be described as: $\mathbf{w}\leftarrow\mathbf{w}-\eta\nabla L(\mathbf{w},b,\mathbf{x}_{i},y_{i})$; $b\leftarrow b-\eta\frac{\partial L(\mathbf{w},b,\mathbf{x}_{i},y_{i})}{\partial b}$ where, $\eta$ is the learning rate, $\nabla L$ is the gradient of the loss function.
[…]
@@
[…]
@@
where, $P(C_{k}|\mathbf{x})$ is the posterior probability of class $C_{k}$, $P(C_{k})$ is the prior probability of class $C_{k}$, $P(\mathbf{x}|C_{k})$ is the likelihood of $\mathbf{x}$ given $C_{k}$, $P(\mathbf{x})$ is the marginal likelihood of $\mathbf{x}$. The likelihood is estimated based on the frequency of features, and parameters for MNB are estimated using Laplace smoothing.
[…]
@@
where, $\mu_{C_{k},i}$ is the mean of $x_{i}$ in $C_{k}$, $\sigma_{C_{k},i}^{2}$ is the variance of $x_{i}$ in $C_{k}$. Further, parameters are estimated using maximum likelihood estimation in GNB.
[…]
Table 4: Hyperparameter Settings For Baseline M2 Models